how do i find the answer to this problem. like what are the steps i need.
3(x-2) = 18
12(2-x)= 6
9/2 (x+3) = 27
thanks

Answer by Isabellaksmith(3)

#1
3 ( x-2 ) = 18
3x-6=18
3x -6 + 6 =18+6
3x = 24
3x/3 = 24/3
x = 8
plug in x ( 8 )
3 ( 8-2 ) = 18
3 ( 6 ) =18
3 * 6 = 18
18=18

#2
12 ( 2-x ) = 6
24 - 12x = 6
24-24 -12x =6 -24
-12x = -18
-12x/-12=-18/-18
x = -18/-12
#3 9/2 (x+3) = 27
9/2 times 2/9 ( x+3 ) = 27 times 2/9
( x+3 ) =6
x+3-3 = 6-3
x = 3
plug in
9/2 ( 3+3) =27
9/2 times 6 =27
27 = 27